OpenStudy (anonymous):

Determine wheather each equation represents a direct variatation. If it does, find the costant of variation. 2x+3y=0 someone help me please?

jimthompson5910 (jim_thompson5910):

solve for y, what do you get?

OpenStudy (anonymous):

like do 3divided by 0??

jimthompson5910 (jim_thompson5910):

no, you first subtract 2x from both sides

jimthompson5910 (jim_thompson5910):

then you divide both sides by 3

OpenStudy (anonymous):

you get 3 because 2- 3 =1 1devided by 3 = 3

jimthompson5910 (jim_thompson5910):

2x+3y=0 3y=0-2x 3y = -2x y = -2/3x

jimthompson5910 (jim_thompson5910):

the equation is now in the form y = kx where k = -2/3

OpenStudy (anonymous):

i see

jimthompson5910 (jim_thompson5910):

so that means that this is an equation of direct variation

OpenStudy (anonymous):

ohh okay i get it now